two charged particle q1=10uC and q2=20uC are separated by distance r=15cm.what is the magnitude of the electrostatic force with solution.


1
Expert's answer
2022-03-09T10:32:46-0500

The Coulomb's law says


"F=k\frac{q_1q_2}{r^2}""F_1=9*10^9*\frac{10*10^{-6}*20*10^{-6}}{(0.15)^2}=80\:\rm N"